OpenStudy (anonymous):

Express in simplest form: x2-1/x2+3x+2 Helpp please

OpenStudy (lgbasallote):

\[x^2 - 1 = (x+1)(x-1)\] \[x^2 + 3x + 2 = (x+1)(x+2)\] therefore \[\frac{x^2-1}{x^2 + 3x + 2} = \frac{(x+1)(x-1)}{(x+1)(x+2)}\]

OpenStudy (lgbasallote):

cancel like terms in numerator and denominator

OpenStudy (lgbasallote):

do you get that @taty555 ?

OpenStudy (anonymous):

ohh I understand now

OpenStudy (lgbasallote):

great!!